# On Mean Field Games in Infinite Dimension

### v1

@@ 7. Examples from Stochastic Optimal Control
Typically, the Hamiltonian $\mathcal{H}$ comes from stochastic optimal control problems. Consider the following $H$-valued stochastic optimal control problem for which the value function
$v(t,x)=\inf_{\alpha\in\mathcal{A}_{t}}\mathsf{E}\left[\int_{t}^{T}f(X^{t,x,\alpha(\cdot)}_{s},\alpha_{s},m(s))\mathrm{d}s+g(X^{t,x,\alpha(\cdot)}_{T},m(T))\right],$
@@
$\mathcal{A}_{t}=\big{\{}\alpha:[t,T]\times\Omega\rightarrow\Lambda\ \ (\mathcal{F}_{s}^{t})-\mbox{progressively measurable}\big{\}}$
for some reference probability space $(\Omega,\mathcal{F},\mathcal{F}_{s}^{t},\mathsf{P},W)$, $\Lambda$ is a complete separable metric space, $m(\cdot)\in\mathcal{S}$, and $X^{t,x,\alpha(\cdot)}$ solves, in mild sense, on $[t,T]$ the infinite dimensional SDE
$dX_{s}=(AX_{s}+b(X_{s},\alpha_{s},m(s)))\mathrm{d}s+\mathrm{d}W_{s},\ \ \ \ X_{t}=x.$
@@
$\mathcal{H}(x,p,\mu)=\sup_{\alpha\in\Lambda}\big{\{}-b(x,\alpha,\mu)p-f(x,\alpha,\mu)\big{\}}.$
@@
    $b(x,\alpha,\mu)=-\alpha+b_{0}(x)$, where $b_{0}:H\rightarrow\mathbb{R}$ is bounded and Lipschitz continuous;
@@
    $f(x,\alpha,\mu)=f_{0}(x,\mu)+f_{1}(|\alpha|)$, where $f_{0}:H\times\mathcal{P}_{1}(H)$ is continuous in all variables, $f_{0}(\cdot,\delta_{0})$ is bounded, $f_{0}$ is Lipschitz continuous in $m$ with respect to ${\bf d}_{1}$, uniformly in $x$, and $f_{1}\in C^{1,1}_{\rm loc}(\mathbb{R})$ is symmetric and uniformly convex [^5] This means that there exists $\eta>0$ such that $s\mapsto f_{1}(s)-\eta s^{2}$ is convex.;
@@
$\mathcal{H}(x,p,\mu)=\sup_{|\alpha|\leq R}\big{\{}\langle\alpha,p\rangle-f_{1}(|\alpha|)\big{\}}-\langle b_{0}(x),p\rangle-f_{0}(x,\mu)=:\mathcal{H}^{1}(p)-\langle b_{0}(x),p\rangle-f_{0}(x,\mu)$
and in this case
$\mathcal{H}^{0}(x,p)=\mathcal{H}^{1}(p)-\langle b_{0}(x),p\rangle,\quad F(x,\mu)=f_{0}(x,\mu),\quad G(x,\mu)=g(x,\mu).$
It is easy to see that Assumption 2.6(i) is satisfied. Regarding Assumption 2.6(ii), we compute explicitly that
@@
Since $f_{1}\in C^{1,1}(\mathbb{R})$ and is uniformly convex, we deduce that $(f_{1}^{\prime})^{-1}$ is Lipschitz continuous, so $H$ is Lipschitz continuous in $p$ too. Moreover, we have
@@
and it is clear that this function is Lipschitz continuous. Hence, $\mathcal{H}^{0}_{p}(x,p)=D\mathcal{H}^{1}(p)-b_{0}(x)$ is bounded and Lipschitz continuous in $p$, uniformly in $x\in H$.
Examples of functions satisfying Assumption 6.1(iii) when $H=\mathbb{R}^{d}$ are for instance in Section 3.4.2 of [10], volume I, and similar examples also work in a real separable Hilbert space $H$ and satisfy the conditions imposed in this paper. We present two examples.
@@
$\int_{H}[F_{i}(x,\mu_{1})-F_{i}(x,\mu_{2})](\mu_{1}-\mu_{2})(\mathrm{d}x)=\left|\int_{H}h_{i}(y)m(\mathrm{d}y)\right|^{2},\quad i=1,2.$
@@
Let $\ell:H\times[0,\infty)\rightarrow\mathbb{R}$ be bounded, continuous, and such that $\ell(x,\cdot)$ is strictly increasing for every $x\in H$ and is Lipschitz continuous with Lipschitz constant independent of $x$. Let $\rho:H\rightarrow[0,\infty)$ be bounded and Lipschitz continuous. Let $\nu$ be a positive finite measure on $H$ which is full  [^6] That is $\nu(D)>0$ for every open set $D$ in $H$. An example of a measure which is full is the nondegenerate Gaussian measure with mean $a\in H$ and covariance operator $Q\in\mathcal{L}_{1}^{+}(H)$, see [13, Proposition 1.25] (nondegeneracy means that ${\rm ker}\,Q=\{0\}$).. We define
@@ Appendix A Compactness in the space of probability measures
Let $\{\mu_{n}\}\subseteq S$. By [36, Ch. VI, Th. 2.2] [^7] In [36, Ch. VI, Th. 2.2] it is stated that the second condition alone guarantees relative compactness of $S$. This is clearly not true, as can be seen by simply taking as $S$ a non compact sequence concentrated on the line generated by $\mathbf{e}_{1}$. However the proof of this theorem there remains valid if we add the assumption of uniform boundedness of second moments., the set $S$ is tight. Hence, by Prokhorov’s theorem, it is weakly relatively compact in $\mathcal{P}_{1}(H)$; that is, there exists $\mu\in\mathcal{P}_{1}(H)$ and a subsequence, still labeled by $\{\mu_{n}\}$, such that
$\int_{H}\phi(x)\mu_{n}(\mathrm{d}x)\rightarrow\int_{H}\phi(x)\mu(\mathrm{d}x),\ \ \ \forall\phi\in C_{b}(H).$
The claim follows using Lemma A.1 and [38, Th. 7.12(ii)].
@@ Appendix B Classical and strong solutions to Kolmogorov equations
In this section, we define classical, mild, and $\mathcal{K}$-strong solutions of backward Kolmogorov equations and recall an approximation result used in Section 6. The material is mainly taken from [19, Section B.7] and is presented, for the reader’s convenience, in a simplified form needed here. We first consider the following terminal value problem (Kolmogorov equation) in the Hilbert space $H$:
@@
We call the function $u$ above, defined by the right hand side, the mild solution of (B.1). Recalling the definition of $L_{0}$ in (3.3), we define on $D(L_{0})$ the norm
@@
Arguing as in Theorem 2.7 of [16], it can be proved that $D(L_{0})$, endowed with the above norm, is a Banach space [^8] The definition of the domain of the operator studied in this paper is slightly different, but the arguments of the mentioned reference can be adapted easily. We also notice that the Banach space structure is not essential for our purposes, even if it simplifies the notation.. We now recall a notion of classical solution to (B.1) (see [19, Definition B.82]) [^9] We also observe that there are other definitions of classical solutions in the literature, see e.g. Section 6.2 of [17, Section 6.2], [25, Definition 4.6], [26, Definition 4.1]. We refer interested readers to [19, Section B.7.1] for a detailed overview on that..
@@
To introduce a notion of a strong solution, called $\mathcal{K}-$strong solution, we first recall a special case of the definition of $\mathcal{K}-$convergence which is needed in this paper (see [19, Definition B.84]).
@@
The definition of a $\mathcal{K}$-strong solution to (B.1) provided below is a special case of a more general definition that can be found in [19, Definition B.85].
@@
We end this section with an approximation result which is a stronger version of [19, Theorem B.95(i)].
@@
We can apply [19, Theorem B.95 (i)] (in the case when, in the notation there, $m=0$, $\eta(t)=t^{-\gamma}$), to conclude that $u$ is also a $\mathcal{K}$-strong solution of the same equation. Moreover, from the proof of [19, Theorem B.95 (i)] one can easily see that the approximating data $\varphi_{n}$, $f_{n}$, and consequently the approximating solutions $u_{n}$, can be chosen so that $\varphi_{n}\in D(L_{0})$ and $f_{n},\partial_{t}\bar{v}_{n}\in C_{b}([0,T]\times H)$ for all $n\in\mathbb{N}$. As a consequence, we get that $u_{n}\in\mathcal{D}_{T}$ for all $n\in\mathbb{N}$.

### v2

@@ 7. Examples from Stochastic Optimal Control
A typical application of our framework is when the Hamiltonian $\mathcal{H}$ arises from stochastic optimal control problems. To illustrate that, consider the $H$-valued stochastic optimal control problem with value function
(7.1) $v(t,x)=\inf_{\alpha_{\cdot}\in\mathcal{A}_{t}}\mathsf{E}\left[\int_{t}^{T}f(X^{t,x,\alpha_{\cdot}}_{s},\alpha_{s},m(s))\mathrm{d}s+g(X^{t,x,\alpha_{\cdot}}_{T},m(T))\right],$
@@
$\mathcal{A}_{t}=\big{\{}\alpha_{\cdot}:[t,T]\times\Omega\rightarrow\Lambda\ \ (\mathcal{F}_{s}^{t})-\mbox{progressively measurable}\big{\}},$
defined on some reference probability space $(\Omega,\mathcal{F},(\mathcal{F}_{s}^{t})_{s\geq t},\mathsf{P},W)$, where $\Lambda$ is a complete separable metric space, $m(\cdot)\in\mathcal{S}$, and $X^{t,x,\alpha_{\cdot}}$ solves, in mild sense on $[t,T]$, the infinite dimensional SDE
(7.2) $\mathrm{d}X_{s}=(AX_{s}+b(X_{s},\alpha_{s},m(s)))\mathrm{d}s+\mathrm{d}W_{s},\ \ \ \ X_{t}=x.$
@@
$\mathcal{H}(x,p,\mu)=\sup_{\alpha\in\Lambda}\big{\{}-\langle b(x,\alpha,\mu),p\rangle-f(x,\alpha,\mu)\big{\}}.$
@@
    $b(x,\alpha,\mu)=-\alpha+b_{0}(x)$, where $b_{0}:H\rightarrow{H}$ is bounded and Lipschitz continuous;
@@
    $f(x,\alpha,\mu)=f_{0}(x,\mu)+f_{1}(|\alpha|)$, where $f_{0}:H\times\mathcal{P}_{1}(H)\rightarrow\mathbb{R}$ is continuous in all variables, $f_{0}(\cdot,\delta_{0})$ is bounded, $f_{0}$ is Lipschitz continuous in $m$ with respect to ${\bf d}_{1}$, uniformly in $x$, and $f_{1}\in C^{1,1}_{\rm loc}(\mathbb{R})$ is symmetric and uniformly convex [^4] This means that there exists $\eta>0$ such that $s\mapsto f_{1}(s)-\eta s^{2}$ is convex.;
@@
$\mathcal{H}(x,p,\mu)$ $=\sup_{|\alpha|\leq R}\big{\{}\langle\alpha,p\rangle-f_{1}(|\alpha|)\big{\}}-\langle b_{0}(x),p\rangle-f_{0}(x,\mu)$
$=:\mathcal{H}^{1}(p)-\langle b_{0}(x),p\rangle-f_{0}(x,\mu).$
In this case
$\mathcal{H}^{0}(x,p)=\mathcal{H}^{1}(p)-\langle b_{0}(x),p\rangle,\quad F(x,\mu)=f_{0}(x,\mu),\quad G(x,\mu)=g(x,\mu),$
and it is easy to see that Assumption 2.6(i) is satisfied. Regarding Assumption 2.6(ii), we compute explicitly
@@
Since $f_{1}\in C^{1,1}(\mathbb{R})$ and is uniformly convex, we deduce that $(f_{1}^{\prime})^{-1}$ is Lipschitz continuous, so $\mathcal{H}$ is Lipschitz continuous in $p$ too. Moreover, we have
@@
and clearly this function is Lipschitz continuous. Hence, $\mathcal{H}^{0}_{p}(x,p)=D\mathcal{H}^{1}(p)-b_{0}(x)$ is bounded and Lipschitz continuous in $p$, uniformly in $x\in H$. This makes our existence result, Theorem 5.6, applicable.
As for uniqueness, examples of functions $F$ satisfying Assumption 6.1(iii) when $H=\mathbb{R}^{d}$ are for instance in Section 3.4.2 of [12], volume I, and similar examples also work in a real separable Hilbert space $H$ and satisfy the conditions imposed in this paper. We present two examples.
@@
$\int_{H}[F_{i}(x,\mu_{1})-F_{i}(x,\mu_{2})](\mu_{1}-\mu_{2})(\mathrm{d}x)=\left|\int_{H}h_{i}(y)(\mu_{1}-\mu_{2})(\mathrm{d}y)\right|^{2},\quad i=1,2.$
@@
Let $\ell:H\times[0,\infty)\rightarrow\mathbb{R}$ be bounded, continuous, and such that $\ell(x,\cdot)$ is strictly increasing for every $x\in H$ and is Lipschitz continuous with Lipschitz constant independent of $x$. Let $\rho:H\rightarrow[0,\infty)$ be bounded and Lipschitz continuous. Let $\nu$ be a positive finite measure on $H$ which is full  [^5] That is $\nu(D)>0$ for every open set $D$ in $H$. An example of a measure which is full is the nondegenerate Gaussian measure with mean $a\in H$ and covariance operator $Q\in\mathcal{L}_{1}^{+}(H)$, see [15, Proposition 1.25] (nondegeneracy means that ${\rm ker}\,Q=\{0\}$).. We define
@@
7.1. An economic model
We illustrate here an application to an economic model in the presence of capital heterogeneity. Inspired by the model presented in [1, Sec. 3(a)], a mean field game model of competition among firms operating in the same sector was studied in [9]. Precisely, in [9], the production capacity of a representative firm is assumed to follow the dynamics
$\mathrm{d}K_{s}=-\delta K_{s}\mathrm{d}s+\sigma K_{s}\mathrm{d}B_{s}+I_{s}\mathrm{d}{s},\ \ K_{0}=k>0,$
where $\delta,\sigma>0$, $(I_{s})_{s\geq 0}$ is a nonnegative real valued process representing the investment rate of the firm, and $B$ is a standard one dimensional Brownian motion. According to the discussion in [1, Sec. 3(a)], the goal of a representative firm is to solve
$v(t,k):=\inf_{I_{\cdot}}\mathsf{E}\left[\int_{t}^{T}\left(-\pi(K_{s},\mathcal{K}(m(s)))+\frac{1}{2}I_{s}^{2}\right)\mathrm{d}s\right],$
where $I_{\cdot}$ belong to a suitable set of admissible control processes,
$\mathcal{K}(\mu)=\left(\int_{\mathbb{R}}k^{\theta}\mu(\mathrm{d}k)\right)^{1/\theta}\ \ \ \ \ \ \mu\in\mathcal{P}_{\theta}(\mathbb{R}),$
with $\theta>0$ and $\pi:(0,\infty)^{2}\rightarrow\mathbb{R}$, $(k,\mathcal{K})\mapsto\pi(k,\mathcal{K})$ is nondecreasing in $k$ and nonincreasing in $\mathcal{K}$ [^6] In [9], the case $\theta=1$ and $\pi(k,\mathcal{K})=\mathcal{K}^{-\beta}k$, where $\beta>0$, is considered.
This model can be modified as follows. First of all, in order to fit our setting, we assume that the noise is additive777 This is done to fit our additive noise setting. Note that, in this way, the capacity may become negative and this is of course an unpleasant fact from the point of view of applications. It would thus be interesting to investigate, in future works, cases with non-additive noise which may ensure that the infinite dimensional process $X$ lies in the positive cone of $H=L^{2}(S^{1};\mathbb{R})$; see, e.g., [30]. with the diffusion coefficient $1$; that is, we look at the dynamics of the form
$\mathrm{d}K_{s}=-\delta K_{s}\mathrm{d}s+\mathrm{d}B_{s}+I_{s}\mathrm{d}{s}.$
We may then enrich the model assuming that, rather than having a large number of single firms, we have a large number of networks of firms, each network being distributed around a circle $S^{1}$. The location of the firm in the representative network is parametrized by $a\in S^{1}$ and capital $K$ depends on this parameter. Therefore, the representative network’s state is represented by a parametrized stochastic process $K_{s}(a)$, where $s\in[t,T]$ and $a\in S^{1}$. We may assume, following the modeling of [8], that also the investment rate process $I_{s}(a)$ is parametrized by the location $a\in S^{1}$, and that capital $K_{s}(a)$ naturally evolves in time-space according to a depreciation law and, by taking account of the “spatial” dimension $a\in S^{1}$, according to a diffusion law; that is, according to the overall law
$\frac{\partial^{2}K_{s}(a)}{\partial a^{2}}-\delta K_{s}(a).$
Assuming also that the noise depends also on the location $a$, this leads to the following controlled SPDE:
(7.3) $\mathrm{d}K_{s}(a)=\left(\frac{\partial^{2}K_{s}(a)}{\partial a^{2}}-\delta K_{s}(a)\right)\mathrm{d}s+\mathrm{d}B_{s}(a)+I_{s}(a)\mathrm{d}s,\ \ \ K_{0}(a)=k(a).$
We consider the following generalization of the optimization problem above:
(7.4) $v(t,k(\cdot))=\inf_{I_{\cdot}(\cdot)}\mathsf{E}\left[\int_{t}^{T}\left(\int_{S^{1}}\left(-\pi(K_{s}(a),\mathcal{K}(m(s)))+\frac{1}{2}I_{s}(a)^{2}\right)\mathrm{d}a\right)\mathrm{d}s\right],$
where
(7.5) $\mathcal{K}(\mu)=\int_{H}\left(\int_{S_{1}}f(k(a))\mathrm{d}a\right)\mu(\mathrm{d}k),\ \ \ \mu\in\mathcal{P}_{1}(H),$
where $f:\mathbb{R}\rightarrow\mathbb{R}$ is a given Lipschitz and bounded function. Formally setting $W_{s}=B_{s}(\cdot)$ and assuming that $W$ is a cylindrical Brownian motion in the Hilbert space $H$, and formally setting $x=k(\cdot)$, $X_{s}=K_{s}(\cdot)$, $\alpha_{s}=-I_{s}(\cdot)$, the SPDE (7.3) can be viewed in the setting of the present section as a controlled SDE (7.2) in the Hilbert space $H=L^{2}(S^{1};\mathbb{R})$ with the following specifications:
$A=\frac{\partial^{2}}{\partial a^{2}}-\delta,\ \ \ b(x,\alpha)=-\alpha.$
Similarly, the optimal control problem can be reformulated as (7.1) by setting $g=0$ and
$f(x,\alpha,\mu)=\frac{|\alpha|^{2}}{2}+F(x,\mu),$
where
$F(x,\mu)=-\int_{S^{1}}\pi(x(a),\mathcal{K}(\mu))\mathrm{d}a.$
Imposing the constraint $I_{t}(\cdot)\in\Lambda=B_{R}$ and assuming that $\pi$ is Lipschitz and bounded, the conditions for our existence result (Theorem 5.6) are fulfilled.
Regarding uniqueness, we assume that $\pi(k,\mathcal{K})=\pi_{1}(\mathcal{K})f(k)$, where $\pi_{1}$ is Lipschitz continuous, nonnegative, bounded, and strictly decreasing, and $f$ is the same function appearing in (7.5) and required now to be Lispschitz continuous, nonnegative, nondecreasing, and bounded. Then, for every $\mu_{1},\mu_{2}\in\mathcal{P}_{1}(H)$, we have
$\int_{H}[F(x,\mu_{1})-F(x,\mu_{2})](\mu_{1}-\mu_{2})(\mathrm{d}x)$
$=$ $-\int_{H}\left[\int_{S^{1}}\big{(}\pi(x(a),\mathcal{K}(\mu_{1}))-\pi(x(a),\mathcal{K}(\mu_{2}))\big{)}\mathrm{d}a\right](\mu_{1}-\mu_{2})(\mathrm{d}x)$
$=$ $-\big{(}\pi_{1}(\mathcal{K}(\mu_{1}))-\pi_{1}(\mathcal{K}(\mu_{2}))\big{)}\int_{H}\left(\int_{S^{1}}f(x(a))\mathrm{d}a\right)(\mu_{1}-\mu_{2})(\mathrm{d}x)$
$=$ $-\big{(}\pi_{1}(\mathcal{K}(\mu_{1}))-\pi_{1}(\mathcal{K}(\mu_{2}))\big{)}(\mathcal{K}(\mu_{1})-\mathcal{K}(\mu_{2}))\geq 0.$
with equality if and only if $\mathcal{K}(\mu_{1})=\mathcal{K}(\mu_{2})$. This shows that:
  1. (i)
    Assumption 6.1(iii) is fulfilled;
  2. (ii)
    Since
    $\big{(}\pi_{1}(\mathcal{K}(\mu_{1}))-\pi_{1}(\mathcal{K}(\mu_{2}))\big{)}(\mathcal{K}(\mu_{1})-\mathcal{K}(\mu_{2}))=0\ \Longrightarrow\ \mathcal{K}(\mu_{1})=\mathcal{K}(\mu_{2}),$
    also Assumption 6.1(iv)(b) holds.
The remaining requirements of Assumption 6.1 are fulfilled by construction. Hence, our uniqueness result, Theorem 6.2, applies.
@@ Appendix A Compactness in the space of probability measures
By [39, Ch. VI, Th. 2.2] [^8] In [39, Ch. VI, Th. 2.2] it is stated that the second condition alone guarantees relative compactness of $S$. This is clearly not true, as can be seen by simply taking as $S$ a non compact sequence concentrated on the line generated by $\mathbf{e}_{1}$. However the proof of this theorem there remains valid if we add the assumption of uniform boundedness of second moments., the set $S$ is tight. The claim follows using [2, Prop. 7.1.5].
@@ Appendix B Classical and strong solutions to Kolmogorov equations
In this section, we define classical, mild, and $\mathcal{K}$-strong solutions of backward Kolmogorov equations and recall an approximation result used in Section 6. The material is mainly taken from [21, Section B.7] and is presented, for the reader’s convenience, in a simplified form needed here. We first consider the following terminal value problem (Kolmogorov equation) in the Hilbert space $H$:
@@
We call the function $u$ above, defined by the right hand side, the mild solution of (B.1). Recalling the definition of $L_{0}$ in (3.3), we define
(B.3) $D(L_{0})$ =$=\Big{\{}\phi\in UC^{2}_{b}(H):\ D^{2}\phi\in UC_{b}(H;\mathcal{L}_{1}(H)),\ A^{*}D\phi\in UC_{b}(H;H)\Big{\}}.$
and endow $D(L_{0})$ with the norm
@@
Arguing as in Theorem 2.7 of [18], it can be proved that $D(L_{0})$, endowed with the above norm, is a Banach space [^9] The definition of the domain of the operator studied in this paper is slightly different, but the arguments of the mentioned reference can be adapted easily. We also notice that the Banach space structure is not essential for our purposes, even if it simplifies the notation.. We now recall a notion of classical solution to (B.1) (see [21, Definition B.82]) [^10] We also observe that there are other definitions of classical solutions in the literature, see e.g. Section 6.2 of [19, Section 6.2], [27, Definition 4.6], [28, Definition 4.1]. We refer interested readers to [21, Section B.7.1] for a detailed overview on that..
@@
To introduce a notion of a strong solution, called $\mathcal{K}-$strong solution, we first recall a special case of the definition of $\mathcal{K}-$convergence which is needed in this paper (see [21, Definition B.84]).
@@
The definition of a $\mathcal{K}$-strong solution to (B.1) provided below is a special case of a more general definition that can be found in [21, Definition B.85].
@@
We end this section with an approximation result which is a stronger version of [21, Theorem B.95(i)].
@@
We can apply [21, Theorem B.95 (i)] (in the case when, in the notation there, $m=0$, $\eta(t)=t^{-\gamma}$), to conclude that $u$ is also a $\mathcal{K}$-strong solution of the same equation. Moreover, from the proof of [21, Theorem B.95 (i)] one can easily see that the approximating data $\varphi_{n}$, $f_{n}$, and consequently the approximating solutions $u_{n}$, can be chosen so that $\varphi_{n}\in D(L_{0})$ and $f_{n},\partial_{t}\bar{v}_{n}\in C_{b}([0,T]\times H)$ for all $n\in\mathbb{N}$. As a consequence, we get that $u_{n}\in\mathcal{D}_{T}$ for all $n\in\mathbb{N}$.